Which of the following statements are True and
which are False ? Justify your answer with a
short proof or by a counter-example.

(a) The operation *, defined by x * y = log (xy) is
a binary operation on S, where
S={xER x>0}.

(b) If a and b are eigenvalues of two n x n
matrices A and B respectively, then a + b is
an eigenvalue of A + B.

(c) If S and T are linear transformations such
that SoT is defined and is 1 — 1, then S is
1 — 1.

(d) T : R³- R³: T((x1, x2, x3), (y 1 , y2, y3)) =
(x1 + x2 + x3) . (y1 + y2 + y3)
is an inner product on R³.

(e) {India, — 5, Jamila} is a set.
1
Expert's answer
2020-12-08T07:38:54-0500

(a) The operation "*", defined by "x * y = \log (xy)" is not a binary operation on

"S=\{x\in \mathbb R \ |\ x>0\}," bacause "1*1=\log(1\cdot 1)=\log 1=0\notin S."


Answer: false


(b) Let us consider two "2\times 2" matrices "A=\left[\begin{array}{cc} 1 & 1 \\ 1 & 1 \end{array}\right]" and "B=\left[\begin{array}{cc} -1 & 0 \\ 3 & -1 \end{array}\right]". Let us find their eigenvalues:

"\det(A-aE)=\left|\begin{array}{cc} 1-a & 1 \\ 1 & 1-a \end{array}\right|=(1-a)^2-1=0" implies "1-a=\pm 1", and therefore, "a=0" or "a=2;"

"\det(B-bE)=\left[\begin{array}{cc} -1-b & 0 \\ 3 & -1-b \end{array}\right]=(-1-b)^2=0" implies "b=-1;"


"\det(A+B-cE)=\left|\begin{array}{cc} -c & 1 \\ 4 & -c \end{array}\right|=c^2-4=0" implies "c=2" or "c=-2".


It follows that "a=0" and "b=-1" are the eigenvalues of the matrices "A" and "B" respectively, but "a+b=-1" is not the eigenvalue of the matrix "A+B."


Answer: false


(c) Let define the linear transformations "T:\mathbb R^2\to\mathbb R^4,\ \ T(x_1,x_2)=(x_1, x_2, 0, 0)" and "S:\mathbb R^4\to\mathbb R^2,\ \ S(x_1,x_2,x_3,x_4)=(x_1, x_2)". Then "S\circ T:\mathbb R^2\to\mathbb R^2." Since "S\circ T(x_1,x_2)=S(x_1, x_2, 0, 0)=(x_1,x_2)", "S\circ T" is an identity map of "\mathbb R^2", and thus is 1 — 1. On the other hand, "S(1,1,0,0)=(1, 1)=S(1,1,1,1)", and consequently, "S" is not 1 — 1.


Answer: false


(d) Let "T :\mathbb R^3\to\mathbb R^3,\ \ T((x_1, x_2, x_3), (y_1 , y_2, y_3)) =(x_1 + x_2 + x_3)\cdot(y_1 + y_2 + y_3)"


Let "u=(x_1,x_2,x_3), \ v=(y_1,y_2,y_3)," and "w=(z_1,z_2,z_3)" be vectors and "\alpha"  be a scalar, then:

1. "T(u+v, w) =((x_1+y_1) + (x_2+y_2) + (x_3+y_3))\cdot(z_1 + z_2 + z_3)=(x_1+ x_2+ x_3+y_1+y_2 +y_3)\cdot(z_1 + z_2 + z_3)=(x_1+ x_2+ x_3)\cdot(z_1 + z_2 + z_3)+(y_1+y_2 +y_3)\cdot(z_1 + z_2 + z_3)=T(u, w)+T(v, w)"


2. "T(\alpha u, v)=T(\alpha(x_1, x_2, x_3), (y_1 , y_2, y_3)) =(\alpha x_1 + \alpha x_2 + \alpha x_3)\cdot(y_1 + y_2 + y_3)=\alpha (x_1 + x_2 + x_3)\cdot(y_1 + y_2 + y_3)=\alpha T( u, v)"


3. "T( u, v) =(x_1 + x_2 + x_3)\cdot(y_1 + y_2 + y_3)=(y_1 + y_2 + y_3)\cdot (x_1 + x_2 + x_3)=T( v, u)"


4.  "T( u, u) =(x_1 + x_2 + x_3)\cdot(x_1 + x_2 + x_3)=(x_1 + x_2 + x_3)^2\ge0". On the other hand, for "(-1,1,0)\ne(0,0,0)" we have "T((-1,1,0),(-1,1,0))=(-1+1+0)(-1+1+0)=0".


Therefore, "T((x_1, x_2, x_3), (y_1 , y_2, y_3)) =(x_1 + x_2 + x_3)\cdot(y_1 + y_2 + y_3)" is not an inner product on "\mathbb R^3".


Answer: false


(e) {India, — 5, Jamila} is a set with three elements: India, — 5 and Jamila.


Answer: true